OpenStudy (anonymous):

simplify (2√5+3√7)^2 show steps

OpenStudy (anonymous):

pls i really need help

OpenStudy (anonymous):

on simplifying just like (a+b)^2 = a^2+2ab +b^2 we get 40+2*2*3*\[\sqrt{35}\]+63 =103+6*square root of (35)

OpenStudy (anonymous):

\[(2\sqrt{5}+3\sqrt{7})(2\sqrt{5}+3\sqrt{7})\]

OpenStudy (anonymous):

so lost o.o

OpenStudy (anonymous):

\[4\sqrt{25}+6\sqrt{35}+6\sqrt{35}+9\sqrt{49}\]

OpenStudy (anonymous):

could you explain in like steps?

OpenStudy (anonymous):

When it is squared, that means that you multiply it by itself, like \[4^{2}= 4\times4\]

OpenStudy (anonymous):

Ok,(2\sqrt{5}+3\sqrt{7})(2\sqrt{5}+3\sqrt{7})

OpenStudy (anonymous):

Oh,oops

OpenStudy (anonymous):

That looks a little confusing

OpenStudy (anonymous):

very lol

OpenStudy (anonymous):

If you do not get it, I will try to explain

OpenStudy (anonymous):

how does it help me simplify

OpenStudy (anonymous):

It will show you how to simplify

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

(a+b)^2=a^2+2*a*b +b^2

OpenStudy (anonymous):

\[(2\sqrt{5} + 3\sqrt{7})^{2}\] \[(2\sqrt{5} + 3\sqrt{7}) (2\sqrt{5} + 3\sqrt{7})\] Because \[(a+b)^{2} = a^{2} + 2(a)(b) + b^{2}\] Therefore\[(2\sqrt{5} + 3\sqrt{7}) (2\sqrt{5} + 3\sqrt{7}) = 20 + 6\sqrt{35} + 63\] Therefore \[6\sqrt{35} +83\]

OpenStudy (anonymous):

omg that's all i wanted thank you sooo much XD

OpenStudy (anonymous):

Your welcome :)

OpenStudy (anonymous):

@PinkPanda me??

OpenStudy (anonymous):

yesh

OpenStudy (anonymous):

@PinkPanda No problem ^.^

OpenStudy (anonymous):

wait did you mean 83 or 43?

OpenStudy (anonymous):

I meant 83

OpenStudy (anonymous):

so you subtracted 20 from each side?

OpenStudy (anonymous):

@BABYShawol184

OpenStudy (anonymous):

No I multiplied the two binomials which gave me the answer \[20 + 6\sqrt{35} +83\] Then I simplified by adding 20 and 83 which gave me 83

OpenStudy (anonymous):

oh ok lol

OpenStudy (anonymous):

Does that help??? :S